OpenStudy (anonymous):

Fun Land theme park has a total of 80 males and females who are employed at the park. If the number of males is 20 more than twice the number of females, how many males are employed at the park?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

do u want me to explain

OpenStudy (anonymous):

no its okay thank you

OpenStudy (anonymous):

could you help me with another?

OpenStudy (anonymous):

@cdr123 you are supposed to explain the process to them not just give an answer

OpenStudy (anonymous):

i know you asked to explain but maybe next time start with an explanation rather than the answer

OpenStudy (anonymous):

ok tousii45

OpenStudy (anonymous):

sorry if i sounded mean :/

OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8

OpenStudy (anonymous):

which equation would you like to use? @allie_bear22

OpenStudy (anonymous):

i think the first 1

OpenStudy (anonymous):

do you want the options of answers?

OpenStudy (anonymous):

@touseii45

OpenStudy (anonymous):

no its fine we can find the correct answer then compare them to the list you have(:

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay for the first one you have 4x + y = 11 subtract 4x from each side so you can use the this equation as substitution in the second equation

OpenStudy (anonymous):

x+y=7

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

i did wat you said

OpenStudy (anonymous):

@cdr123 do u have the answer cuz this is taking forever

OpenStudy (anonymous):

no you were supposed to subtract 4x from each side not just 4

OpenStudy (anonymous):

6361.7inches2

OpenStudy (anonymous):

wait wat @cdr123 thats not even an option lol

OpenStudy (anonymous):

@cdr123 !! if an ambassador or a moderator were to see you do that you would be in trouble

OpenStudy (anonymous):

@electrokid maybe u can help me

OpenStudy (anonymous):

wait wat did u 2 come up with

OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8

OpenStudy (anonymous):

(3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (anonymous):

those are the options

OpenStudy (anonymous):

whitch one is it?

OpenStudy (anonymous):

from first equation \[y=11-4x\] substitute thi sin the second \[x+2(11-4x)=8\] solve for "x" in this one

OpenStudy (anonymous):

i dont k now how

OpenStudy (anonymous):

thats why i was helping you @allie_bear22 read the code of conduct and you will understand that we cant just give you answers

OpenStudy (anonymous):

expand the second one first

OpenStudy (anonymous):

in the second equation distribute the 2

OpenStudy (anonymous):

i know but i need someone to do the work so i can see wat to do thats how i learn how to do things @touseii45

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

\[x+2(11)-2(4x)=8\\ x+22-8x=8\]find "x"

OpenStudy (anonymous):

how do u find x?!?!?!?

OpenStudy (anonymous):

@touseii45 please continue from here.. I have to leave

OpenStudy (anonymous):

i dont know what to do i learn diffrent from everone else

OpenStudy (anonymous):

@cdr123 can you do the work and not give me the anser cuz its hard for me

OpenStudy (anonymous):

okay i will help you with the work but i wont give you answers

OpenStudy (anonymous):

i thought the anser is d

OpenStudy (anonymous):